OpenStudy (kingethan8500):

The height of a cone-shaped container is 15 centimeters and its radius is 14 centimeters. Kate fills the container completely with honey. Every day Kate uses 40 cubic centimeters of honey from the container to make honey cookies. After how many days will the container be completely empty? Round your answer to the nearest whole number. (Use π = 3.14.) 11 77 82 93

OpenStudy (kingethan8500):

@kg1975

jhonyy9 (jhonyy9):

do you know what is the formula for volume of a cone ?

jhonyy9 (jhonyy9):

V = area of base time height divide 3 this will help you ?

jhonyy9 (jhonyy9):

kingethan do you like collaborate please ? so i will help you fine the right answer

jhonyy9 (jhonyy9):

you find the right answer - sorry

OpenStudy (kingethan8500):

no i did not find it.

jhonyy9 (jhonyy9):

area of base = pi*r^2 area b. = 3,14*14^2 = 615,4 cm^2 V = 615,4 *15 = ? cm^3 so to find in how many days will be empty this container just you need divide the volume of honey from container by this quantity what use Kate every day

jhonyy9 (jhonyy9):

@ILovePuppiesLol do you agree it ? please

jhonyy9 (jhonyy9):

ty

ILovePuppiesLol (ilovepuppieslol):

i am not sure about this @jhonyy9 sorry

jhonyy9 (jhonyy9):

@welshfella do you agree it please ? thank you

OpenStudy (welshfella):

the volume of the cone is (1/3) pi r^2 h = (1/3) * 3.14 *14^2 * 15 = 3077.2 divide this by 40 to get the number of days